# Tree Attention: Topology-Aware Decoding for Long-Context Attention on GPU Clusters

### v3

@@ Abstract
Self-attention is the core mathematical operation of modern transformer architectures and is also a significant computational bottleneck due to its quadratic complexity in the sequence length. In this work, we derive the scalar energy function whose gradient computes the self-attention block, thus elucidating the theoretical underpinnings of self-attention, providing a Bayesian interpretation of the operation and linking it closely with energy-based models such as Hopfield Networks. Our formulation reveals that the reduction across the sequence axis can be efficiently computed in parallel through a tree reduction. Our algorithm, for parallelizing attention computation across multiple GPUs enables cross-device decoding to be performed asymptotically faster (up to $8\times$ faster in our experiments) than alternative approaches such as Ring Attention, while also requiring significantly less communication volume and incurring $2\times$ less peak memory. Our code is publicly available here: https://github.com/Zyphra/tree_attention.
1.  Introduction
*Denotes Equal Contribution
Correspondence to: vasu@zyphra.com, jonathan@zyphra.com
The self-attention operation is the core computational building block of the transformer architecture Bahdanau et al. (2014); Vaswani et al. (2017), which has become an ubiquitous and highly effective workhorse architecture currently applied at scale to language Brown et al. (2020); Kaplan et al. (2020); Hoffmann et al. (2022); Team et al. (2023); Achiam et al. (2023), vision Dosovitskiy et al. (2020), audio Betker (2023), and decision-making Chen et al. (2021); Reed et al. (2022). Nonetheless, the quadratic time complexity of self-attention means that significant resources are required to train and generate from transformer-based Large Language Models (LLMs), especially for models with large context lengths.
@@ 1.  Introduction
To speed up inference and alleviate memory requirements, recent works have attempted to alter the attention mechanism itself, either by linearizing it Katharopoulos et al. (2020), or approximating it by a kernel map Choromanski et al. (2020); Peng et al. (2021); Arora et al. (2024), which reduces the complexity to linear at the cost of reduced expressiveness. Others have invented alternative sequence mixing architectures such as state-space models which are designed to be efficiently computable in linear time and constant memory Gu and Dao (2023); Dao and Gu (2024); Katsch (2023); Sun et al. (2023); Glorioso et al. (2024). Other approaches utilize efficient algorithms to reduce the computational burden of attention while keeping the core computation the same. These include memory-efficient attention Rabe and Staats (2021), Flash Attention Dao et al. (2022) and Flash Decoding fla (2024), which provide a set of IO-aware kernels to map the attention operation to the GPU hardware resources in an extremely efficient way, significantly reducing the memory overhead required. Further works Character AI (2024); Kang et al. (2024); Liu et al. (2024); Nawrot et al. (2024) explore compressing or otherwise reducing the KV cache required in generation. Finally, Ring Attention Liu et al. (2023) proposes a way to parallelize the attention computation across the sequence axis between GPUs, thus enabling significantly longer contexts than can be served on a single GPU. This is the regime of primary interest of this paper. By leveraging the exact energy function for the self-attention block, we develop a method to speed up inference for long context use-cases when keys and values are sharded across multiple GPUs along the sequence axis.
Our proposed algorithm for computing attention via the gradient of the energy function is built on top of an efficient parallel computation and tree reduction communication strategy. In particular, this formulation lets us devise an asymptotically faster algorithm for performing decoding in which the number of communication steps scales logarithmically with the number of devices, instead of linearly in alternatives such as Ring Attention Liu et al. (2023). Our topology-aware approach illustrated in Fig. 1. significantly outperforms leading attention parallelization methods such as Ring Attention on multiple devices.
@@
In this work, we make three core contributions:
  • We provide a mathematical form for the energy function of self-attention.
  • From this theory, we develop an algorithm for parallelizing the attention computation across devices, leveraging tree-reduction topology.
  • Finally, we show empirically that our algorithm provides asymptotic speedups when decoding across many GPUs on long sequence lengths, as well as large reductions in communication volume and peak memory usage.
2.  Related Work
Following the ubiquitous success of the transformer architecture, there has been significant effort to mathematically understand the nature and meaning of the attention operation and link it to related architectures, such as recurrent networks and Hopfield Networks. A number of recent works have attempted to study self-attention mathematically through the lens of energy functions. Ramsauer et al. (2020) pioneered this field by performing a similar but distinct analysis to relate self-attention with the modern Hopfield networks Krotov and Hopfield (2016), providing a novel and insightful interpretation of self-attention as performing hetero-associative memory lookups using a high-powered nonlinear similarity function Krotov (2021); Millidge et al. (2022). This work was later extended by Hoover et al. (2023), who derived a modified version of the transformer based off an energy function. Beyond this, other recent work attempted a Bayesian reformulation of attention by deriving a probabilistic generative model which matches the operations performed in a self-attention operation Singh and Buckley (2023).
However, while it has long been known that the softmax operation can be derived as the gradient of the following scalar function:
known as the log-sum-exp, an equivalent function for the self-attention block has not yet been derived.
To advance beyond this, our energy function depends on an auxiliary source vector $\zeta$ (see section 4.). Taking the gradient with respect to $\zeta$ yields the exact self-attention operation, as we show below. Consequently, this formulation enables us to express self-attention with no weight tying.
Another notable related work is Feng et al. (2024), where the authors made similar observations as we do in section 5. about how the associative operations within the attention computation can be efficiently parallelized to motivate an attention-based modified RNN architecture for sequence modeling.
3.  Self-Attention
@@
$z^{a}=\sum^{N}_{i=1}\text{softmax}(q_{a}\cdot k_{i}^{T})v_{i}\,.$
The dot product, more explicitly, reads:
$q_{a}\cdot k_{i}=\sum^{d}_{A=1}q_{a,A}k_{i,A},\forall A\in\{1,\cdots,d_{h}\}.$
The capital Latin indices ($A$) span the hidden dimension $d_{h}$. Naively computing attention in this way requires materializing the $qk$ matrix with computational and memory cost quadratic in the sequence length.
Memory-efficient attention Rabe and Staats (2021) is an iterative way to compute the softmax similarities without ever having to materialize the full attention matrix. It performs the following operations, one query (or a chunk of queries) at a time:
@@
Flash attention Dao et al. (2022) utilizes a similar approach to reduce the memory and computational cost of attention, but the algorithm is not adapted for multi-GPU computation. Flash attention performs the iterative algorithm of Rabe and Staats (2021) in a blockwise manner, utilizing the block-parallel computational primitives available inside single GPU tensor cores. Additionally, it precisely sizes the blocks such that they can fit into the SRAM of the GPU for the entire attention computation, effectively performing kernel fusion and preventing many unnecessary IO operations.
4.  Self-Attention as a Gradient

### v4

@@ Abstract
Self-attention is the core mathematical operation of modern transformer architectures and is also a significant computational bottleneck due to its quadratic complexity in the sequence length. In this work, we derive the scalar energy function whose gradient computes the self-attention block, thus elucidating the theoretical underpinnings of self-attention. Our formulation reveals that the reduction across the sequence axis can be efficiently computed in parallel through a tree reduction. Our algorithm, called Tree Attention, for parallelizing exact attention computation across multiple GPUs enables cross-device decoding to be performed asymptotically faster (up to $8\times$ faster in our experiments) than state-of-the-art approaches such as Ring Attention, while also requiring significantly less communication volume and incurring $2\times$ less peak memory. We demonstrate that Tree Attention speeds up decoding up to 4x on Llama 3.1-8B and can be applied to a variety of hardware and networking setups such as H100 DGX nodes, AMD MI300x nodes, and PCIe connected NVIDIA RTX 4090s. Our code is publicly available here: https://github.com/Zyphra/tree_attention
1 Introduction
The self-attention operation is the core computational building block of the transformer architecture Bahdanau et al. (2014); Vaswani et al. (2017), which has become an ubiquitous and highly effective workhorse architecture currently applied at scale to language Brown et al. (2020); Kaplan et al. (2020); Hoffmann et al. (2022); Team et al. (2023); Achiam et al. (2023); Pilault et al. (2023), vision Dosovitskiy et al. (2020), audio Betker (2023), and decision-making Chen et al. (2021); Reed et al. (2022). Nonetheless, the quadratic time complexity of self-attention means that significant resources are required to train and generate from transformer-based Large Language Models (LLMs), especially for models with large context lengths.
@@ 1 Introduction
To speed up inference and alleviate memory requirements, recent works have attempted to alter the attention mechanism itself, either by linearizing it Katharopoulos et al. (2020), or approximating it by a kernel map Choromanski et al. (2020b); Peng et al. (2021); Arora et al. (2024), which reduces the complexity to linear at the cost of reduced expressiveness. Others have invented alternative sequence mixing architectures such as state-space models which are designed to be efficiently computable in linear time and constant memory Gu & Dao (2023); Dao & Gu (2024); Katsch (2023); Sun et al. (2023); Glorioso et al. (2024).
@@
Other approaches utilize efficient algorithms to reduce the computational burden of attention while keeping the core computation the same. These include memory-efficient attention Rabe & Staats (2021), Flash Attention Dao et al. (2022) and Flash Decoding fla (2024), which provide a set of IO-aware kernels to map the attention operation to the GPU hardware resources in an extremely efficient way, significantly reducing the memory overhead required. Further works Character AI (2024); Kang et al. (2024); Liu et al. (2024); Nawrot et al. (2024) explore compressing or otherwise reducing the KV cache required in generation. Finally, Ring Attention Liu et al. (2023) proposes a way to parallelize the attention computation across the sequence axis between GPUs, thus enabling significantly longer contexts than can be served on a single GPU. Since our proposed method is an exact calculation of attention [^1] It can be shown empirically that Ring Attention and Tree Attention are exact computations of Attention since both methods have exactly the same activations as the forward pass of Vanilla Attention., it is a plugin replacement for any multi-GPU sequence parallel mechanism such as the state of the art Ring Attention mechanisms. By leveraging the exact energy function for the self-attention block, we develop a method to speed up inference for long context use-cases when keys and values are sharded across multiple GPUs along the sequence axis.
Our proposed algorithm for computing attention via the gradient of the energy function is built on top of an efficient parallel computation and tree reduction communication strategy. In particular, this formulation lets us devise an asymptotically faster algorithm for performing decoding in which the number of communication steps scales logarithmically with the number of devices, instead of linearly in alternatives such as Ring Attention Liu et al. (2023). Our topology-aware approach illustrated in Fig. 1 significantly outperforms leading attention parallelization methods such as Ring Attention on multiple devices.
2 Related Works
The computational complexity of self-attention, introduced by Vaswani et al. (2017), poses challenges for long sequences due to its quadratic dependency on sequence length, $O(n^{2}\cdot d)$. To address this, attention approximation mechanisms like Linformer (Wang et al., 2020) and Performer (Choromanski et al., 2020a) reduce complexity to linear $O(n)$ using low-rank projections and kernelized approximations on a single device. Sparse models such as Longformer (Beltagy et al., 2020) and BigBird (Zaheer et al., 2020) further optimize computations by restricting attention to local windows or sparsity patterns, significantly reducing resource demands while maintaining performance for specific tasks. Such methods however provide approximations to the attention mechanism while we seek to parallelize the exact attention computation across the sequence axis.
Theoretical work has also contributed to improving the efficiency of both exact and approximate methods. Kernel-based approaches, such as those by Tsai et al. (2019), suggest alternative formulations to self-attention that are computationally efficient. Surveys like Tay et al. (2020) highlight these advancements, emphasizing the synergy between parallelization strategies and sparsity or approximation techniques, ensuring self-attention remains scalable even in constrained computational environments. It must be noted as well that Duman Keles et al. (2022) established lower bounds on the computational complexity of self-attention, demonstrating that achieving sub-quadratic time complexity is unlikely unless the Strong Exponential Time Hypothesis (SETH) is false.
In addition to approximation methods, several approaches focus on parallelizing the exact attention computations. FlashAttention (Dao et al., 2022), for instance, reorganizes the attention computation into smaller, memory-efficient blocks that leverage GPU memory hierarchies to enable faster and parallelized processing of exact attention, and by doing so reduces the memory complexity from quadratic to linear. Other techniques use optimized matrix operations and tiling strategies to distribute attention computations across cores or threads efficiently (Shen et al., 2021). While these methods aim to maximize throughput while maintaining the precision of exact attention, they focus on speeding up single-device attention computation. Since we parallelize exact attention across multiple devices, Ring Attention (Liu et al., 2023) is most comparable to our work. Finally, to the best of our knowledge, there are no other techniques that explore multi-device parallel decoding as we have done in this paper.
3 Self-Attention
@@
$z^{a}=\sum^{N}_{i=1}\mathrm{softmax}(q_{a}\cdot k_{i}^{T})v_{i}\,.$
Naively computing attention in this way requires materializing the $qk$ matrix with computational and memory cost quadratic in the sequence length. Memory-efficient attention Rabe & Staats (2021) is an iterative way to compute the softmax similarities without ever having to materialize the full attention matrix. It performs the following operations, one query (or a chunk of queries) at a time:
@@
Flash Attention Dao et al. (2022) utilizes a similar approach to reduce the memory and computational cost of attention, but the algorithm is not adapted for multi-GPU computation. Flash Attention performs the iterative algorithm of Rabe & Staats (2021) in a blockwise manner, utilizing the block-parallel computational primitives available inside single GPU tensor cores. Additionally, it precisely sizes the blocks such that they can fit into the SRAM of the GPU for the entire attention computation, effectively performing kernel fusion and preventing many unnecessary IO operations.
4 Self-Attention as the Gradient of an Energy Function
Following the ubiquitous success of the transformer architecture, there has been significant effort to mathematically understand the nature and meaning of the attention operation and link it to energy models (Krotov & Hopfield, 2016; Krotov, 2021; Millidge et al., 2022; Hoover et al., 2024), such as Hopfield Networks (Ramsauer et al., 2020; D’Amico & Negri, 2024). Ramsauer et al. (2020) pioneered this field by performing a similar but distinct analysis to relate self-attention with the modern Hopfield networks, providing a novel and insightful interpretation of self-attention as performing hetero-associative memory lookups using a high-powered nonlinear similarity function. This work was later extended by Hoover et al. (2023), who derived a modified version of the transformer based off an energy function. However, while it has long been known that the softmax operation can be derived as the gradient of the following scalar function:
known as the log-sum-exp, an equivalent function for the self-attention block has not yet been derived. We develop in this paper a link between attention and energy functions by introducing an auxiliary source vector $\zeta$, which represents the “external contributions” to the system’s energy (Hopfield, 1982). The source $\zeta$ is the parameter with respect to which we compute the gradient of the scalar energy function to obtain the self-attention operation. As we will see, we need the source in order to write down the generating function of the moments of the distribution since taking the gradient with respect to $\zeta$ yields the exact self-attention operation.
This insight allows us to make the following observation:
Observation 1.
Attention can be expressed at the gradient of an scalar energy function $F(\zeta)$ with respect to the source $\zeta$, such that:
$\sum^{N}_{a=1}\mathrm{softmax}(q\cdot k_{a})v_{a}=\frac{\partial F}{\partial\zeta}\bigg|_{\zeta=0},$ (5)
where the moment generating function (i.e. the energy function) $F(\zeta)$ is defined as:
$F(\zeta)=\log\sum_{a}\exp\left(q\cdot k_{a}^{T}+\zeta\cdot v^{T}_{a}\right).$ (6)
The proof of Observation 1 can be found in Appendix C.1. Please note that this formulation also allows to make a Bayesian interpretation of Attention in Appendix C.2 and motivates our Tree Attention algorithm in the next Section 5.
5 Tree Attention
5.1 Efficient Energy Function Computation
$F_{dec}=\log\sum^{N}_{a=1}\exp(q\cdot k_{a}^{T}+\zeta\cdot v^{T}_{a})\equiv\textrm{logsumexp}_{a}(\{q\cdot k^{T}_{a}+\zeta\cdot v^{T}_{a},a=1,\cdots,N\}).$ (7)
$\textrm{logsumexp}_{a}(\{T_{a},\textrm{logsumexp}_{a}(\{R_{a},S_{a}\})\})=\textrm{logsumexp}_{a}(\{\textrm{logsumexp}_{a}(\{T_{a},R_{a}\}),S_{a}\}),$
$\max_{a}(\{\max_{a}(\{T_{a},R_{a}\}),S_{a}\})=\max_{a}(\{T_{a},\max_{a}(\{R_{a},S_{a}\})\}).$
Theorem 1.
The proof of Theorem 1 is in Appendix E.
[…]
Algorithm 1 Single Query Energy Forward (calculating logsumexp)
[…]
7:Save $lse,m$ for gradient w.r.t $\zeta$.
8:Return lselse
5.2 Efficient parallel decoding
In our case, we want to compute the gradient of the energy function with respect to $\zeta_{A}$ and then set it to zero. This can be done with automatic differentiation engines having set $\zeta$ to be a tensor of zeros from the very outset. We can however manually implement a gradient with respect to $\zeta$ pass of the above Algorithm  1 that does not materialize $\zeta$ in Algorithm 2 below. Note in particular that when we set $\zeta_{A}=0,\,\,A\in\{1,\cdots,d_{h}\}$ then lselse involves only the logsumexp of the dot product between queries and keys.
Algorithm 2 Tree Decoding (using atomic operation on each device)
[…]
2:Calculate $m$ and lselse using Algorithm  1.
3:Scatter a copy of $\mathbf{q},m$ and lselse, and each $\mathbf{k}_{\hat{a}},\mathbf{v}_{\hat{a}}$ to each of the $p$ processors.
4:In parallel compute $r_{\hat{a}}=\mathbf{q}\cdot\mathbf{k}^{T}_{\hat{a}}-m$
5:Compute $R_{\hat{a}}=\frac{\exp(r_{\hat{a}})}{\exp(lse)}\cdot v_{\hat{a}}=\exp(r_{\hat{a}}-lse)\cdot v_{\hat{a}}$
[…]
7:Return $z$
Notice here that by storing $lse,m$ for the backward pass, the only remaining reduction operation that needs to be performed is the one in line 5 of the above algorithm. This single reduction takes $O(N/p)$ time to compute the local sums on each device and $\log p$ time to communicate and combine partial results, and therefore we get the same asymptotic complexity as the logsumexp calculation.
Algorithm 3 Tree Decoding (using Flash Attention 2 on each device)
[…]
2:Use Flash Attention 2 to compute $\textrm{o}=\frac{\sum_{\hat{a}}\exp(\mathbf{q}\cdot\mathbf{k}_{\hat{a}}^{T})\mathbf{v}_{\hat{a}}}{\sum_{\hat{b}}\exp(\mathbf{q}\cdot\mathbf{k}^{T}_{\hat{b}})}$ and $\textrm{lse}=\log\sum_{\hat{b}}\exp(\mathbf{q}\cdot\mathbf{k}^{T}_{\hat{b}})$.
3:Recompute the global max $m=\textrm{{Allreduce}}(\textrm{max},\textrm{lse})$.
4:Get local numerator and denominator by computing: $\textrm{n}=\textrm{o}*\exp(\textrm{lse}-m),\textrm{d}=\exp(\textrm{lse}-m)$.
5:Compute global numerator and denominator with: $\textrm{n}_{g}=\textrm{{Allreduce}}(\textrm{sum},\textrm{n}),\textrm{d}_{g}=\textrm{{Allreduce}}(\textrm{sum},\textrm{d})$.
6:Return result $z=\frac{\textrm{n}_{g}}{\textrm{d}_{g}}$.
5.3 Efficient collective operations using topology-awareness
Communication overheads
While the theoretical analysis above indicates that we should see speedups when using tree-based reductions, this is not necessarily guaranteed in practice due to various potential overheads. In particular, our argument for the time complexity of our proposed Tree Decoding algorithm assumes that communication of partial results is instantaneous, which in practice is never the case. In fact, as we scale the sequence length, or the number of GPUs especially to the multi-node setting, the time taken for communication is the dominant contribution to the total execution time. However, importantly, beyond its asymptotic benefits, Tree Attention benefits from taking advantage of the two-level topology which is standard in modern GPU clusters.
Figure 2: NCCL Send/Recv between two H100 GPUs intra-node and inter-node. GPU clusters offer a two-tier topology where intra-node bandwidth is significantly higher than inter-node. Algorithms such as Tree Attention exploit this topology by reducing inter-node communication requirements, enabling better overlap of communication with computation.
We benchmark our algorithm against a previously proposed sequence parallel attention algorithm called Ring Attention. Like our algorithm, Ring Attention assumes that the sequence is sharded across GPUs and performs the attention computation without gathering all of the sequence on to a single device. Instead, it communicates shards of the keys and values in a point-to-point manner between neighboring GPUs that are logically arranged in a ring topology. This communication is overlapped with the computation of the local shard of the output. In contrast with this strategy, our algorithm scatters the query and communicates the partial result across all GPUs when performing the AllReduce operation, but does not move the key and value shards between GPUs. Consequently, in the decoding case, our method benefits from having lower communication volume and suffers less from the communication cost overhead than Ring Attention does.
Implications of network bandwidth heirarchy
Ring Attention is inherently not topology-aware, and only scales within a network of homogeneous bandwidth. However, this is in conflict with the two-level network topology of modern GPU clusters, which use high-bandwidth interconnects within nodes (NVLINK or PCIe) and comparatively lower-bandwidth interconnects across nodes (InfiniBand or Ethernet). The interconnects greatly differ in bandwidth and latency (see Figure 2). Therefore, Ring Attention is bottlenecked by the slowest interconnect, and cannot always overlap the attention computation with communication. We discuss this point further in 6.3 Tree Attention improves on Ring Attention by using network topology-aware communication patterns to increase overlap of computation and communication, and decrease this scalability bottleneck on communication from the distributed attention computation.
In our empirical experiments , we use Flash Attention 2 (Dao, 2023) within each device, both for our algorithm and for Ring Attention [^2] A JAX-based Ring Attention implementation that uses Flash Attention 2 can be found here: https://github.com/nshepperd/flash_attn_jax.. We provide a simple JAX implementation of our method in Appendix D. Note that our method mirrors Flash Decoding (fla, 2024) except in that case, the parallelization happens at the level of different streaming multiprocessors (SMs) within a GPU whereas we parallelize between different GPUs. All computations are performed in BF16.
6 Results
Similar to Ring Attention, Tree Attention is an exact computation of attention. Since training and evaluation metrics are the same as for attention, our experimental results are focused primarily on latency in section 6.1, peak memory usage in section 6.2 and communication volumes in section 6.3. Since our algorithm computes numerically identical results as the forward pass of standard attention, our performance results transfer seamlessly to transformer architectures.
We performed experiments in Sections 6.1 to 6.3 on a DGX H100 cluster consisting of 16 nodes, each containing 8 H100 GPUs. All GPUs within the node are connected via an all-to-all NVLINK 4.0 (900GBps) topology. Nodes are connected to each other via 8 InfiniBand NDR interconnects per node (1 per GPU), each of which provides 400 Gbps (leading to an aggregate 3.2 Tbps node injection bandwidth).
We also show Ring Attention and Tree Attention comparisons when used in a Llama 3 model (Grattafiori et al., 2024) in Sections 6.4 and C.3 on viarous GPU and interconnect types: 8 H100 GPUs with NVLINK 4.0, 8 AMD MI300X GPUs with AMD infinity fabric for intra-node communication and RoCE for inter-node communication, and 2 RTX 4090 GPUs with PCIe interconnect.
6.1 Latency
Our latency results shows how Tree Attention improves over Ring Attention as we increase the sequence length in Fig. LABEL:fig:relative_times and increase the number of GPUs in Fig. LABEL:fig:absolute_times. To better highlight execution time trends with an increasing sequence length, we have also added relative execution time of both methods with respect to the execution time of ring attention at a sequence length of 80k. With relative execution time in Fig. LABEL:fig:relative_times, we notice that Tree attention’s execution time flattens as the number of GPUs increases, while Ring Attention relative execution time continues to increase. As the plots demonstrate, as we scale the sequence length or the number of GPUs, the gap between Tree Attention and Ring Attention execution time widens asymptotically. Remarkably, Tree Attention achieves close $\times 8$ speedups when we use 128 GPUs on a sequence length of 5.12M. We expect this trend to continue for larger sequence lengths. Please note that our DGX cluster is made up of 16 nodes each with 8 GPUs. Results for 8 GPUS use one node, for 64 GPUs uses 8 nodes and for 128 GPUs uses 16 nodes.
6.2 Memory cost
(b) Peak memory usage of a single attention block with Tree Attention vs Ring Attention when sharded between two RTX 4090s. Results were taken using the JAX memory profiler on one GPU. The difference in peak memory scales with hidden size and sequence length.
[…]
$\textrm{Mem}_{\text{ring}}$ $=4btd+2bd$ (8)
$\textrm{Mem}_{\text{tree}}$ $=2btd+2bd+2bn_{h},$ (9)
6.3 Communication volume
For Ring Attention’s P2P communication strategy, the total volume of data being communicated between devices (in units of number of tensor elements) per iteration scales with $p$ and is given by:
$V_{ring}=2btd\times p$ (10)
$V_{\texttt{Allreduce}}=2\times\frac{p-1}{p}\times\textrm{numel}.$ (12)
$\textrm{numel}\left(\textrm{n},\textrm{d},m\right)=bd+2bn_{h}.$ (13)
Note that we first perform on device the local reductions to obtain the local numerator and denominator on each device which consequently makes it so that $t$, i.e. the size of the local sequence chunk does not appear in the above expression. We then obtain:
$V_{Tree}=2\frac{p-1}{p}\times\left(bd+2bn_{h}\right).$ (14)
Our theoretical analysis shows that per iteration our algorithm maintains a lower communication volume than Ring Attention. Note however that Ring Attention when performed in the training setting with many queries overlaps communication and computation so as to hide its communication costs. However, overlapping communication and computation in the decoding case is infeasible because of how fast the attention computation on a single GPU is relative to how long it takes to communicate the chunk of keys and values between two devices.
Concretely, let us take the example of decoding from a context of length 640000 split between $8$ GPUs within one node. Let us take a hidden size of 2048 and fix our data type to be bfloat16. Each device for decoding takes $O(10^{-5})$ seconds to perform the Flash Attention computation. The time it takes to move the keys and values of the corresponding size between adjacent GPUs as per Fig. 2 is roughly $O(10^{-3})$ seconds. The latency incurred between nodes is even greater and therefore overlapping is not feasible due to this disparity in timescales.
6.4 Performance with a LLama Transformer Model
To show that Tree attention can also be used in real world applications, we also measured end-to-end throughput with the Llama 3.1 8B model Grattafiori et al. (2024) on prompt sequences of length 32k, 64k, 128k and 256k using ring attention or tree attention for decoding (with prefill) 10 tokens in Table 1. We ran these experiments on 8 H100 GPUs in a DGX cluster (connected with NVLink) as well as 4 MI300X GPUs in an AMD cluster connected with AMD infinity fabric. In Table 2 of Appendix C.3, we also show similar throughput results on 2 RTX 4090 GPUs connected with PCIe. In all cases we see that Tree attention for decoding has significantly lower latency than Ring Attention for decoding with a prefill stage. Ring Attention is up to $\times$4 faster using 8x H100s and up to $\times$3 faster using 4x MI300x. We expect this gap to increase as we increase the number of nodes.
While we have previously discussed that Ring Attention works best when used with the Ring Topology of TPU clusters, Table 1 and 2 show that Tree Attention results generalize well to various types of systems, number of GPUs, communication protocols and network topologies.
Table 1: Average Decoding Time (in seconds) with a prefill stage comparisons, using the 8B Llama 3.1 model with Tree Attention (ours) and Ring Attention (SOTA) across various sequence lengths and GPU types. Average results and standard error ($\pm$) are computed using 10 trial runs.
Sequence 8x H100s 4x MI300x
Length Tree Attn Ring Attn Speedup Tree Attn Ring Attn Speedup
32k 0.60 $\pm$ 0.15 2.57 $\pm$ 0.35 $\times$4 1.05 $\pm$ 0.01 3.57 $\pm$ 0.25 $\times$3
64k 1.08 $\pm$ 0.10 4.42 $\pm$ 0.38 $\times$4 2.36 $\pm$ 0.01 7.33 $\pm$ 0.25 $\times$3
128k 2.68 $\pm$ 0.28 6.38 $\pm$ 0.58 $\times$2 6.43 $\pm$ 0.25 16.40 $\pm$ 0.40 $\times$3
256k 2.89 $\pm$ 0.62 8.19 $\pm$ 1.07 $\times$3 15.30 $\pm$ 4.93 35.12 $\pm$ 5.02 $\times$2
7 Discussion and Conclusion
In this paper, we have derived the energy function for self-attention and demonstrated how the computation of the derivative of this function provides a novel and efficient method for computing attention in parallel. This advantage is especially apparent when performing decoding across multiple devices, in which case our Tree Attention enables us to substantially outperform SOTA Ring Attention with an asymptotically superior algorithm, with $\times 8$ speedups when we use 128 GPUs on a sequence length of 5.12M. We also see that the AllReduce operation that we use involves sending partially reduced objects, which greatly reduces the volume of communicated data as well as the peak memory requirement. In a real-world application, using the Llama 3.1 model with 1B and 8B parameters, we find that decoding with a prefill stage using Tree Attention gets us $\times$3-5 speedupds compared to Ring Attention. Further, by testing our method on various types of GPUs clusters including AMD MI300xs, we show that Tree Attention generalizes very well to various communication protocols and network topologies.
Appendix A More Related Work
Recent work has attempted a Bayesian reformulation of attention by deriving a probabilistic generative model which matches the operations performed in a self-attention operation Singh & Buckley (2023). This follows on from a line of work which relates self-attention to the well-studied Hopfield Network architecture Ramsauer et al. (2020). The idea being that while the Hopfield Network is defined in terms of dynamics upon an energy landscape, this same picture can be cast into a Bayesian interpretation by identifying the energy function with a variational free energy functional and hence deriving the generative model that is implicit in the self-attention update.
Another notable related work is Feng et al. (2024), where the authors made similar observations as we do in section 5 about how the associative operations within the attention computation can be efficiently parallelized to motivate an attention-based modified RNN architecture for sequence modeling.
Appendix B More background on the tree reduction operation
A tree reduction operation is a hierarchical strategy to perform a reduction operation (e.g., sum, product, maximum, minimum) over a set of data elements efficiently, especially in parallel computing. This approach reduces the overall computational complexity and enables efficient utilization of parallel processing resources. Here’s how it works:
  • Divide the problem into smaller tasks: The input data is divided into smaller chunks, and the reduction operation is performed pairwise between adjacent elements in these chunks.
  • Form a tree-like structure: The results from the first level of reductions are themselves reduced pairwise in the next level. This continues until the entire dataset has been reduced to a single result.
  • Iterative or recursive aggregation: The aggregation typically follows a binary tree pattern, but other fan-in numbers (e.g., k-ary trees) can also be used. Each node in the tree represents a partial reduction result, and the root of the tree holds the final result.
Because a tree structure has a logarithmic depth to total number of nodes, a tree reduction can asymptotiacally reduce the number of total steps required to perform an operation when it is possible to aggregate partial results, and additionally is amenable to parallelization since k-ary trees can be defined to match the number of available processors for parallel processing. Additionally, many existing networking topologies such as Nvidia’s NVLINK and Infiniband, due to the natural advantages of tree structures, are designed with such a toplogy meaning that tree operations are natural and efficient to perform.
Appendix C Attention as the Gradient of an Energy Function and Bayesian interpretations
C.1 Proof of Observation 1